OpenStudy (anonymous):

Using 6400 km as the radius of Earth, calculate how high above Earth's surface you would have to be in order to weigh 1/16th of your current weight. Show all work leading to your answer OR describe your solution using 3-4 complete sentences.

OpenStudy (anonymous):

Weight if the Force of Gravity. So you need to find a ratio of forces of gravity (a fraction) that gives you 1/16th. Therefore, there will be 2 weights:\(\ w_1\) and\(\ w_2\). By Newton's Universal Law of Gravitation you have that: \[\ F = G\frac{mM}{r^2_1}\] And the ratio: \[\ \frac{w1}{w_2} = \frac{G\frac{mM}{r^2_1}}{G\frac{mM}{r^2_2}}\] \(\ G,\ m \) and\(\ M\)re the same in both cases, and so they cancel out, giving you with: \[\ \frac{w1}{w_2} = \frac{r^2_1}{r^2_2} = \left(\frac{r_1}{r_2}\right)^2 \] Since it is a relationship to the second power if you ingrease \(\ r_2\) 4 times, the strength will decrease \(\ 4^2 = 16\) times.

OpenStudy (anonymous):

Ok, I am very lost.

OpenStudy (anonymous):

What is the part where you get lost?